# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v11

@@ B.6.1 Daily Wage Distribution by Ages
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASCT is $1.05$ yen, meaning that there is approximately $1.5$ yen difference. [^96] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (ten+ or below ten workers) and engines (use or not). I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.979797This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen. In comparison, the overall weighted average is $3.23$ yen, a difference of only $0.02$ yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged $15$ and older is $1.14$, while the overall average is $1.13$, a difference of only $0.01$ yen. This suggests that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial. [^98] The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. If our interest is in the late 20s to the early 40s, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$. [^99] Strictly, this weight may not be required to be in this range. I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one. The monthly wage equation is then specified as follows:
@@ B.10 Distribution of the Idiosyncratic Shocks
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.101101101I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the daily panels include a set of consecutive household-date cells with no income. These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@ B.12 Risk Preferences
Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.102102102Specifically, Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.103103103The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.104104104This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises. Thirteen heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.105105105Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
@@ C.1 Conceptual Framework
where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.106106106In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.5 Testing Relationship between Volatility of Earnings and Use of Credit Purchases
Column (1) of Table C.3 presents the result for Equation 4.107107107 The estimated coefficient is close to zero and statistically insignificant. A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases. Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications. Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households that use credit for their purchases. This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.108108108Despite this, low-income households might have had more difficulty using credit for purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).

### v12

@@ B.6.1 Daily Wage Distribution by Ages
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASCT is $1.05$ yen, meaning that there is approximately $1.5$ yen difference. [^98] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (ten+ or below ten workers) and engines (use or not). I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.999999This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen. In comparison, the overall weighted average is $3.23$ yen, a difference of only $0.02$ yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged $15$ and older is $1.14$, while the overall average is $1.13$, a difference of only $0.01$ yen. This suggests that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.100100100The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. If our interest is in the late 20s to the early 40s, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.101101101Strictly, this weight may not be required to be in this range. I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one. The monthly wage equation is then specified as follows:
@@ B.10 Distribution of the Idiosyncratic Shocks
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the daily panels include a set of consecutive household-date cells with no income. These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@ B.12 Risk Preferences
Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises. Thirteen heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
@@ C.1 Conceptual Framework
where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
In Table C.2, I analyze the relationship between credit purchases and competitiveness among local retailers using the selling item by scale of business cells reported in Table C.1. In Column (1), I regressed the sales in credit purchases (in percentage points) on the number of local competitors. I then conditioned the business scales in Column (2). The $F$-statistic $p$-value for joint zero-slope hypotheses for these scale dummies is estimated to be $0.990$. In Column (3), I included the selling item dummies instead of the scale dummies. The $F$-statistic $p$-value for joint zero-slope hypotheses is rejected at the conventional level, and the measure of fit is relatively high compared with the other two models reported in Columns (2) and (3). The estimate is $3.420$, suggesting that a one unit increase in the local competitors increases the use of credit purchase by roughly $3.4$ percentage points. Given that the standard deviation of the local competitors is approximately $2.3$, the estimated magnitude corresponding to the one standard deviation increase would be $7.9$ ($3.420\times 2.3$). This is roughly $41$% of the standard deviation of the dependent variable ($19.2$) and thus, is an economically meaningful magnitude.
Table C.2: Use of Credit Purchases and Competitiveness among Local Retailers
Dependent Variable: Sales in Credit Purchases (%)
(1) (2) (3)
Number of Competitors 7.1907.190***7.190 ∗ ∗ ∗ 7.2547.254***7.254 ∗ ∗ ∗ 3.4203.420**3.420 ∗ ∗
$[2.041]$ $[2.145]$ $[1.310]$
Sample mean of the DV $59.4$ $59.4$ $59.4$
Model OLSE OLSE OLSE
Number of observations 12 12 12
Number of retail shops (weight) 326 326 326
Scale FEs No Yes No
Selling item FEs No No Yes
$F$-statistic $p$-value for the FEs 0.990 0.000
R-squared 0.577 0.578 0.986
Notes: The observations used in this table are from the selling item by scale of business cells reported in Table C.1. The dependent variable in Columns (1)–(3) is the percentage of sales under credit purchases. “Number of competitors” indicates the number of peers within a three-cho (approximately 327 meters) radius. The sample mean and standard deviation of the dependent variable are $59.4$ and $19.2$, respectively. The sample mean of the number of competitors is $6.89$ (Std. Dev. = $2.26$). All the regressions are weighted by the number of stores in each selling item by scale of business cell (see third column of Table C.1). The scale of business dummies are added in Column 2, whereas the selling item dummies are included in Column 3. $F$-statistics $p$-values for these dummies are reported in Columns 2 and 3. The standard errors are based on the HC2 estimator proposed by Horn et al. (1975).
@@ C.5 Testing Relationship between Volatility of Earnings and Use of Credit Purchases
Column (1) of Table C.4 presents the result for Equation 4.109109109 The estimated coefficient is close to zero and statistically insignificant. A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases. Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications. Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households that use credit for their purchases. This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.110110110Despite this, low-income households might have had more difficulty using credit for purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).